# The Convex Matching Distance in Multiparameter Persistence

### v1

@@ 7 Conclusions
In this work, we introduced the convex matching distance, a new pseudo-metric to compare functions with values in ${\mathbb{R}}^{2}$. If $\boldsymbol{{\varphi}}$ and $\boldsymbol{{\psi}}$ are two such functions, the convex matching distance is defined by maximising the bottleneck distance between the persistence diagrams of the convex combinations $\boldsymbol{{\varphi}}^{t}$ and $\boldsymbol{{\psi}}^{t}$ of the coordinates of $\boldsymbol{{\varphi}}$ and $\boldsymbol{{\psi}}$. Whereas the classical matching distance requires the computation of a two-parameter family of persistence diagrams (see [CeDFFeFrLa13, Landi2022, FrMoQuTo25]), our metric depends only on the one-parameter family indexed by $t\in[0,1]$. This reduction in dimensionality, combined with the additional regularity of the operator sending $\boldsymbol{{\varphi}}$ to $\boldsymbol{{\varphi}}^{t}$, leads to both conceptual simplifications and computational advantages improving the effective computability of the metric. Moreover, the smoothness preservation property of the proposed operator allows for its incorporation into learning-based frameworks, ensuring stable and well-behaved gradients and enabling the use of standard optimization and training procedures.
@@ Acknowledgements
P.F. was supported by GNSAGA-INdAM, CNIT National Laboratory WiLab, and the WiLab-Huawei Joint Innovation Center. U.F. acknowledges the Italian National Biodiversity Future Center (NBFC) - National Recovery and Resilience Plan (NRRP) funded by the EU-NextGenerationEU (project code CN 00000033) and the Innovation Ecosystem Robotics and AI for Socio-economic Empowerment (RAISE) - National Recovery and Resilience Plan (NRRP) funded by the EU-NextGenerationEU (project code ECS 00000035). E.M.G. was supported by GNAMPA-INdAM and GNAMPA Project CUP E53C22001930001. N.Q. was supported by CNIT National Laboratory WiLab, and the WiLab-Huawei Joint Innovation Center. S.S. aknowledges the GNSAGA-INdAM and the MUR Excellence Department Project MatMod@TOV awarded to the Department of Mathematics, University of Rome Tor Vergata, CUP E83C23000330006. F.T. was funded by the Knut and Alice Wallenberg Foundations and the WASP Postdoctoral Scholarship Program.

### v2

@@ 6 Special values
7 Experiments
Throughout the section, every function $\boldsymbol{{\varphi}}=({\varphi}_{1},{\varphi}_{2})\colon X\to{\mathbb{R}}^{2}$ takes values in $0,1]^{2}$. This prevents scale-dependent artifacts in both the convex combinations and the bottleneck distance computations. Moreover, each point cloud will lie in the unit cube. We remind that $\boldsymbol{a}=(a,1-a)$ and $\boldsymbol{b}=(b,-b)$.
The aim of this section is to evaluate the convex matching distance against the classical matching distance. In practice, both distances are approximated by evaluating the maximum over finite grids. Specifically, we evaluate three scenarios:
  • CMD: approximation of $\operatorname{\mathbf{cmd}}_{k}$ by restricting to a uniform grid of $11$ values $t\in\{0,0.1,\dots,1\}$ in the convex combination $(1-t)\varphi_{1}+t\varphi_{2}$;
  • MD 121: approximation of $d_{\mathrm{match},k}$ by restricting to a fine $11\times 11$ grid with $a$ ranging from $0.05$ to $0.95$ and $b\in\{0,0.1,\dots,1\}$ in the parameter space for $\boldsymbol{{\varphi}}^{*}_{\boldsymbol{a},\boldsymbol{b}},\boldsymbol{\psi}^{*}_{\boldsymbol{a},\boldsymbol{b}}$;
  • MD 10: approximation of $d_{\mathrm{match},k}$ by restricting to a coarser $2\times 5$ grid with $a\in\{0.25,0.75\}$ and $b\in\{0,0.25,0.5,0.75,1\}$.
The goal is to show that CMD provides a reliable alternative for MD 121 (which is approximately $11$ times slower), while outperforming MD 10, whose computational cost is comparable to CMD. We conduct three experiments: (i) a direct evaluation of distance quality on MNIST digits (Section 7.1), (ii) a classification task on synthetic shapes (Section 7.2) and (iii) a classification task on chaotic attractors (Section 7.3). We conclude with a dedicated timing analysis (Section 7.4), showing that the empirical computational cost aligns with the expected theoretical speed-up of CMD over the classical matching distance.
All classification experiments use $k$-nearest neighbors ($k=3$) with bottleneck distance (tolerance $\varepsilon=0.01$). Classification is performed using a most-confident selection strategy across $H_{0}$ and $H_{1}$, and performance is evaluated via 10-fold cross-validation. Code and results are available at https://doi.org/10.5281/zenodo.18644292. Persistent homology computations were carried out using the GUDHI library [39].
7.1 Distance quality on MNIST
The aim of this section is to quantify CMD, MD 121 and MD 10 on MNIST, and assess whether the discriminating power of CMD is comparable to that of MD121, which in turn is higher than that of MD10. To keep the computational cost limited, we selected $100$ samples for each of the 10 classes, yielding $\binom{1000}{2}=499,500$ pairs. Let $X$ denote the $28\times 28$ pixel grid. We define the functions ${\varphi}_{1}(x)=I(x)/255$, where $I(x)\in[0,255]$ is the pixel intensity, and ${\varphi}_{2}$ as the radial distance from the image center, restricted to active pixels ($I(x)>50$) and set to 0 otherwise. The function ${\varphi}_{2}$ is normalised to take values in $[0,1]$. For persistence computation, we construct a cubical complex on the $28\times 28$ pixel grid, using the superlevel set filtration induced either by $\boldsymbol{{\varphi}}^{t}$ (for CMD) or $\boldsymbol{{\varphi}}^{*}_{\boldsymbol{{a}},\boldsymbol{b}}$ (for MD 121 and MD 10).
Figure 4 displays scatterplots comparing the three distances. Each point corresponds to an image pair, with coordinates given by the two distances being compared: for instance, in panel (a), the $x$-coordinate is CMD and the $y$-coordinate is MD 121, both computed in homological dimension 0. Points near the diagonal (red dashed line) indicate image pairs where both distances yield similar values, while points far from the diagonal indicate that one distance dominates over the other. As shown in Figure 4 (a, b), CMD and MD 121 are nearly identical in both homological dimensions. In contrast, the comparison between CMD and MD 10 (Figure 4 (c, d)) favors CMD. As a further verification, the Spearman correlation between CMD and MD 121 is $0.980$ ($H_{0}$) and $0.993$ ($H_{1}$), compared to $0.926$ ($H_{0}$) and $0.955$ ($H_{1}$) for MD 10 vs MD 121. This confirms that CMD agrees more with the fine-grid MD than its coarser version, while maintaining a computational cost comparable to the latter.
(a) MNIST: CMD vs MD 121 for $H_{0}$.
(b) MNIST: CMD vs MD 121 for $H_{1}$.
(c) MNIST: CMD vs MD 10 for $H_{0}$.
(d) MNIST: CMD vs MD 10 for $H_{1}$.
Figure 4: Distance comparison for MNIST. (a,b) CMD vs MD 121 for $H_{0}$ and $H_{1}$. (c,d) CMD vs MD 10 for $H_{0}$ and $H_{1}$. Red dashed line shows $y=x$.
7.2 Classification of synthetic shapes
As a second experiment, we use CMD, MD 121, and MD 10 directly as classifiers in a $k$-nearest neighbors framework. For each test sample, we compute the $k=3$ nearest neighbors in both $H_{0}$ and $H_{1}$ separately. We then select the prediction from the homological dimension with higher confidence, defined as the proportion of the $k$ neighbors voting for the most common class. In case of equal confidence, we use the average distance as a tiebreaker. We generate five classes of 3D point clouds in $0,1]^{3}$: circle, sphere, torus, three clusters, and two concentric circles. We refer to Figure 5 for examples of point clouds for each class, colored by eccentricity. The bifiltration uses $\varphi_{1}$ as the distance to the $15$-th neighbor (codensity), and $\varphi_{2}$ as eccentricity [42], both normalized to $[0,1]$. Persistence is computed via Alpha complex with lower-star extension. Each class contains $40$ samples, and we test across additive Gaussian noise levels $\sigma\in\{0.00,0.03,0.06,0.09,0.12\}$ applied independently to each coordinate of the point cloud, and number of points in the point cloud $n\in\{500,1000,5000\}$. As single-parameter baselines, we compute both codensity lower-star and eccentricity lower-star persistence. Table 1 reports classification accuracies (mean $\pm$ standard deviation) across all combinations of sample size $n$ and noise level $\sigma$. The results are consistent across all configurations: CMD and MD 121 achieve comparable performance throughout, while MD 10 systematically underperforms. The key difference emerges at high sample size and moderate-to-high noise. At $n=5000$, $\sigma=0.06$: CMD achieves $97.5\%$, surpassing MD 121 ($94.2\%$), while MD 10 drops to $80.8\%$. At $\sigma=0.12$: CMD ($78.2\%$) and MD 121 ($81.0\%$) remain comparable, while MD 10 falls to $58.0\%$. Figure 6 visualizes this trend for $n=5000$, showing that CMD maintains robust performance as noise increases, while MD 10 degrades significantly at $\sigma\geq 0.09$. Notably, MD 10 performs comparably to the best single-parameter baseline (codensity lower-star, see Table 2), while CMD and MD 121 substantially outperform all 1-PH methods, confirming that the second filtration carries complementary discriminative information.
Circle
Sphere
Torus
3 Clusters
2 Circles
Figure 5: Examples of synthetic 3D point clouds for each class ($n=5000$, $\sigma=0.0$), colored by eccentricity.
Table 1: Classification accuracy (%) on synthetic shapes.
$n=500$ $n=1000$ $n=5000$
$\sigma$ CMD MD 121 MD 10 CMD MD 121 MD 10 CMD MD 121 MD 10
0.00 $91.8\pm 5.3$ $91.0\pm 6.2$ $91.3\pm 5.6$ $94.5\pm 3.9$ $92.5\pm 6.4$ $3.3\pm 4.1 99$$3\pm 0.8 99.$$8\pm 0.5 99.$$9.0\pm 1.1$
0.03 $5.2\pm 2.8 88$$8.0\pm 4.0$ $84.5\pm 3.4$ $94.5\pm 2.5$ $94.2\pm 1.5$ $3.3\pm 2.7 99$$9.3\pm 0.8$ $98.5\pm 0.9$ $97.2\pm 2.2$
0.06 $76.0\pm 3.3$ $80.7\pm 4.4$ $69.5\pm 5.1$ $89.0\pm 2.4$ $92.5\pm 3.6$ $78.0\pm 7.4$ $97.5\pm 1.9$ $94.2\pm 2.5$ $80.8\pm 4.2$
0.09 $74.0\pm 5.9$ $75.0\pm 5.3$ $0.3\pm 1.9 77$$7.3\pm 6.4$ $82.3\pm 6.3$ $78.8\pm 6.9$ $88.5\pm 2.8$ $89.0\pm 3.2$ $86.3\pm 2.6$
0.12 $69.0\pm 4.8$ $69.3\pm 4.9$ $64.0\pm 5.7$ $69.5\pm 5.2$ $74.0\pm 5.4$ $70.0\pm 5.8$ $78.2\pm 2.8$ $81.0\pm 3.5$ $58.0\pm 5.2$
Figure 6: Classification accuracy vs noise for synthetic shapes at $n=5000$. CMD maintains robust performance across all noise levels, while MD 10 degrades significantly at high noise ($\sigma\geq 0.09$).
Table 2: Single-parameter baseline accuracies (%) on synthetic shapes.
$n=500$ $n=1000$ $n=5000$
$\sigma$ Codensity LS Ecc LS Codensity LS Ecc LS Codensity LS Ecc LS
0.00 $83.7\pm 3.7$ $75.2\pm 5.6$ $93.2\pm 3.4$ $71.3\pm 7.8$ $99.8\pm 0.5$ $78.3\pm 6.1$
0.03 $75.0\pm 6.2$ $24.2\pm 3.7$ $91.8\pm 4.4$ $23.5\pm 4.2$ $98.5\pm 1.6$ $25.0\pm 4.0$
0.06 $77.0\pm 4.0$ $20.5\pm 2.5$ $81.7\pm 3.9$ $22.0\pm 2.9$ $84.3\pm 3.7$ $21.0\pm 4.9$
0.09 $76.7\pm 5.8$ $19.2\pm 3.0$ $78.0\pm 4.6$ $24.8\pm 6.8$ $84.0\pm 3.3$ $28.3\pm 8.9$
0.12 $62.5\pm 4.8$ $21.3\pm 2.2$ $69.3\pm 3.5$ $22.2\pm 6.4$ $64.0\pm 5.9$ $25.3\pm 5.2$
7.3 Classification of chaotic attractors
Lastly, we classify five chaotic dynamical systems: Lorenz, Rössler, Thomas, Sprott, and Four-Wing attractors. Chaotic attractors have been widely studied in the topological data analysis literature [33, 35, 43, 45]. Each trajectory is sampled with $n=500$ points and rescaled to $0,1]^{3}$. The bifiltration uses ${\varphi}_{1}=\sqrt{\alpha}$ (Alpha complex filtration values, clipped at $0.15$ for numerical stability) and ${\varphi}_{2}=$ eccentricity (lower-star extension). Clipping avoids degenerate filtration values arising from quasi-coplanar points. We generate 40 samples per class. We refer to Figure 7 for visual examples of trajectories from each attractor class. The considered approximation of the convex matching distance performs comparably in this dataset to the two approximations of the matching distance: CMD $95.2\pm 1.9\%$, MD 121 $96.5\pm 2.4\%$, and MD 10 $93.8\pm 3.4\%$. In fact, the topological signal is strong enough that even a coarse approximation of the matching distance suffices. However, a distance quality analysis as in Section 7.1 (Figure 8) reveals that CMD and MD 121 agree more closely with each other than either does with MD 10, consistent with our previous findings.
Lorenz
Rössler
Thomas
Sprott
Four-Wing
Figure 7: Examples of attractors trajectories.
(a) Attractors: CMD vs MD 121 for $H_{0}$.
(b) Attractors: CMD vs MD 121 for $H_{1}$.
(c) Attractors: CMD vs MD 10 for $H_{0}$.
(d) Attractors: CMD vs MD 10 for $H_{1}$.
Figure 8: Distance comparison for the attractors dataset. (a,b) CMD vs MD 121 for $H_{0}$ and $H_{1}$. (c,d) CMD vs MD 10 for $H_{0}$ and $H_{1}$. Red dashed line shows $y=x$.
7.4 Computational cost
We evaluate execution time on a subset of MNIST (150 images). The cost of computing the distance is dominated by persistence diagram computations and bottleneck distance evaluations. Table 3 reports timings for the three approximation strategies.
Table 3: Wall-clock time (seconds) for distance matrix computation on MNIST subset (150 images).
Method PD computations Bottleneck distances Total
CMD 1.3s 1.5s 2.8s ($1.0\times$)
MD 121 13.2s 12.2s 25.4s ($9.0\times$)
MD 10 1.2s 1.1s 2.2s ($0.8\times$)
The observed speedups of MD 121 and MD 10 over CMD is lower than the theoretical ratio due to differences in persistence diagram sizes. The matching distance parametrisation produces on average smaller and sparser diagrams, especially for $a,b$ close to 0 and $1$. This reduces the cost per bottleneck distance computation. Nevertheless, CMD achieves nearly identical execution time to MD 10 while providing substantially better approximation quality.
@@ 8 Conclusions
In this work, we introduced the convex matching distance, a new pseudo-metric to compare functions with values in ${\mathbb{R}}^{2}$. If $\boldsymbol{{\varphi}}$ and $\boldsymbol{{\psi}}$ are two such functions, the convex matching distance is defined by maximising the bottleneck distance between the persistence diagrams of the convex combinations $\boldsymbol{{\varphi}}^{t}$ and $\boldsymbol{{\psi}}^{t}$ of the coordinates of $\boldsymbol{{\varphi}}$ and $\boldsymbol{{\psi}}$. Whereas the classical matching distance requires the computation of a two-parameter family of persistence diagrams (see [20, 4, 30]), our metric depends only on the one-parameter family indexed by $t\in[0,1]$. This reduction in dimensionality, combined with the additional regularity of the operator sending $\boldsymbol{{\varphi}}$ to $\boldsymbol{{\varphi}}^{t}$, leads to both conceptual simplifications and computational advantages improving the effective computability of the metric. Moreover, the smoothness preservation property of the proposed operator allows for its incorporation into learning-based frameworks, ensuring stable and well-behaved gradients and enabling the use of standard optimization and training procedures.
@@
We validated the practical effectiveness of the convex matching distance through experiments on MNIST, synthetic 3D shapes under varying noise levels, and chaotic dynamical attractors. Our results demonstrate that the convex matching distance achieves accuracy comparable to the full matching distance computed on a fine $11\times 11$ parameter grid, while requiring approximately $10\times$ fewer persistence diagram computations. Moreover, the convex matching distance substantially outperforms coarser approximations of the matching distance, confirming that the convex parametrization effectively captures the discriminative information of the bifiltration.
@@ Acknowledgements
F.C. was funded by Inria. P.F. was supported by GNSAGA-INdAM, CNIT National Laboratory WiLab, and the WiLab-Huawei Joint Innovation Center. U.F. acknowledges the Italian National Biodiversity Future Center (NBFC) - National Recovery and Resilience Plan (NRRP) funded by the EU-NextGenerationEU (project code CN 00000033) and the Innovation Ecosystem Robotics and AI for Socio-economic Empowerment (RAISE) - National Recovery and Resilience Plan (NRRP) funded by the EU-NextGenerationEU (project code ECS 00000035). E.M.G. was supported by GNAMPA-INdAM and GNAMPA Project CUP E53C22001930001. N.Q. was supported by CNIT National Laboratory WiLab, and the WiLab-Huawei Joint Innovation Center. S.S. acknowledges the GNSAGA-INdAM and the MUR Excellence Department Project MatMod@TOV awarded to the Department of Mathematics, University of Rome Tor Vergata, CUP E83C23000330006. F.T. was funded by the Knut and Alice Wallenberg Foundations and the WASP Postdoctoral Scholarship Program.